OpenStudy (shubhamsrg):

a man is standing at point A . he moves one unit in every step he takes . he can move either in horizontal or vertical direction,even backwards or forwards,left or right.. whats is the probability that he is 1 unit away from A after 9 steps ?

OpenStudy (dumbcow):

wow so many possibilities...there are 4^9 different routes he can take question then is to find how many of those routes end up on one of the 4 spaces 1 unit away from A

OpenStudy (shubhamsrg):

yeah!! this ques is amazing!!

OpenStudy (paxpolaris):

in this question the order in which the moves are made do not matter ..... is it ok to say there are \(\Large 4^9/9!\) routes he can take?

OpenStudy (dumbcow):

hmm no, 9! is bigger than 4^9 so that is impossible yes the order matters when talking about the different routes he can take

OpenStudy (dumbcow):

the number of spots he can end up do not depend on order but im not sure what that number is

OpenStudy (dumbcow):

ok i wrote a program and simulated this a million times and the probability he ends up 1 unit from A is about 24.2%

OpenStudy (dumbcow):

seems kinda high...im gonna double check my work

OpenStudy (shubhamsrg):

ohh..well i dont have the ans to this ques..my sir had asked this in the class..this was some 4-5 weeks ago..then we all forgot about it :P i remembered it yesterday and tried all night but couldnt build up any method.. so,,i dont know the ans..0.242 might be correct..

OpenStudy (paxpolaris):

I think i can figure out the number of possible routes that would bring you back to A after 8 steps

OpenStudy (dumbcow):

no it checks out...so answer is around 25%, may be some error due to random number generator

OpenStudy (dumbcow):

@PaxPolaris , that sounds great

OpenStudy (paxpolaris):

then what ....?

OpenStudy (dumbcow):

weird, the same simulation says that there is probability of 0 he ends up back at A in 9 steps

OpenStudy (paxpolaris):

true

OpenStudy (paxpolaris):

not possible with 9 I said 8

OpenStudy (dumbcow):

oh i see now because its odd you never end up at starting place thats why probability of being close to A is so high then

OpenStudy (shubhamsrg):

didnt get you sir,,how do you arrive at this expression?

OpenStudy (dumbcow):

i am very impressed, i have no idea how you came up with that haha thats why i studied computational math...i'll run the simulations while you work on the theories

OpenStudy (shubhamsrg):

sir i really wanna know how you came up with the soln,,please..

OpenStudy (paxpolaris):

While the order in which you take the steps may matter as far as probability is concerned... It does not matter to where you end up.. If order doesn't matter - you can think that he takes 8 steps that cancel each other out and bring him back to the origin ‘A’, and a Final step that takes him to the endpoint. Let's call this Final deciding step: X. In the other 8 steps - there have to be as many up(u) as down(d), and as many Left(L) as right(R). So the total number of u & L is 4 ...(excluding X which can be anything)

OpenStudy (paxpolaris):

So, I set out to calculate the possibilities when we have 4,3,2,1, or 0 u. When we have 4u and no L: uuuu .... (and 4d) & X is u or d .... {the possible ways to order the 9 steps are:\[\large 2\times {9\choose5}\] + plus if X is L or F:\[\large 2\times {9\choose1}{8\choose4}\] And then muliply rhe whole thing, since the number of ways is the same if we have LLLL uuuu / LLLL : \[\Large 2\times \left[ 2\cdot {9\choose5} + 2\cdot9{8\choose4} \right]\]

OpenStudy (paxpolaris):

*** that should say: And then multiply the whole thing by 2, since the number of ways is the same if we have LLLL

OpenStudy (shubhamsrg):

am getting to it..i should give it an independent try again..thank you so much sir ^_^

OpenStudy (paxpolaris):

then do the same for: uuuL / uLLL \[\Large 2\times \left[ 2\cdot{9\choose4}{5\choose1}{4\choose1}+2\cdot{9\choose3}{6\choose2}{4\choose1} \right]\] and for: uuLL\[\Large4\cdot{9\choose3}{6\choose2}{4\choose2} \] finally add the three up, and divide by 4^9

OpenStudy (shubhamsrg):

hmmn,,to be frank i didnt read your latest post...as i said,,i'd like to try on my own first,,i'll refer to it when i fail ( which i surely will).. hmmn.. and again,,a big thank you ^_^